OpenStudy (buttermequeasy):

The formula for degrees Celsius is: C = 5 over 9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 10 degrees Celsius is.

OpenStudy (buttermequeasy):

?????

OpenStudy (stormfire1):

To solve the equation for F: \[C=\frac{5}{9}(F-32)\]\[\frac{9}{5}C=F-32\]\[\frac{9}{5}C+32=F\] For the second part just plug in 10:\[\frac{9}{5}10+32=F\]\[F = 50 degrees\]